OpenStudy (wesdg1978):

If f(x)=3x and g(x)=x-4, evaluate (fogon)(x)

OpenStudy (anonymous):

\((f\circ g\circ n)(x)~?\) What's \(n\)?

OpenStudy (zzr0ck3r):

i think that should say (f of g of)(x) so just f composite g

OpenStudy (anonymous):

If it is just (fog)(x) ie f of g of x, or f composite g like zzrocker said, then you replace all x's in the f(x) function with the whole g(x) function. Ill get you started, f(x)=3x g(x)=x-4 (fog)(x)= 3 ( g(x) )

OpenStudy (wesdg1978):

The answer is 3x^2-15 , I don't know how to get there. I've tried to combine 3x and x-4 as many ways as possible and I still can't get 3x^2-15.

OpenStudy (anonymous):

do me a favor and confirm that you are giving us the question correctly, and make sure that the answer is for the right question. I believe that we are missing out on information from the question.

OpenStudy (wesdg1978):

Let g(x)=x-4, h(x)-x^2-1, and f(x)=3x. Evaluate the following: (fogon)(x)

OpenStudy (anonymous):

is that n suppose to be an h, or vice versa?

OpenStudy (wesdg1978):

The only thing that I can think of is that the n is supposed to be an h.

OpenStudy (wesdg1978):

Yes, I think so

OpenStudy (wesdg1978):

I caught is as I was typing it out.

OpenStudy (anonymous):

It should be, I just worked it out and it came out to your answer. Try it yourself and let me know if you are having any more trouble with it.

OpenStudy (wesdg1978):

I don't see how to get 3x^2-15. I got 3x^4-12x^3-3x^2+12x, which doesn't reduce to 3x^2-15

OpenStudy (anonymous):

Here, this might help you get started: f(x)=3x g(x)=x-4 h(x)=x^2 -1 Start out with just goh(x) first: goh(x)=(x^2-1)-4 then use that in fogoh(x): fogoh(x)=3((x^2-1)-4)

OpenStudy (anonymous):

note that (x^2-1)-4 is NOT distributive property

OpenStudy (wesdg1978):

Ok, so you get 3(x^2-5) which comes out to the answer, but where did the x in x-4 go?

OpenStudy (anonymous):

it also gets substituted in the equation, the x in x-4 became h(x), or x^2-1. so: g(x)=x-4 goh(x)= (h(x))-4 notice that h(x) substitutes for x goh(x)= (x^2-1)-4 same thing for f(x)

OpenStudy (wesdg1978):

OK, I see it. For some reason I was trying to multiply it all out instead of substituting in place of the x's. Thanks for your help!

OpenStudy (anonymous):

no problem =)